Circular No. 7407

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1118+480 C. A. Haswell, Open University; R. I. Hynes, Southampton University;

and A. R. King, Leicester University, on behalf of a larger collaboration, report that Hubble Space Telescope observations of XTE J1118+480 were made during Apr. 8.52–8.81 UT: “Preliminary reductions reveal a flux density at 150 nm of 3 × 10^{−13} erg cm^{−2} s^{−1} Å^{−1} with a continuum slope from 115 to 800 nm somewhat flatter than the ν^{1/3} power law expected for a viscously heated blackbody accretion disk. This may suggest that the intrinsic x-ray flux is relatively low, removing the motivation for suggesting a nearly edge-on viewing angle (IAUC 7392). The Balmer jump appears in absorption. N V emission (124.0, 124.3 nm) is most prominent, with FWHM 1.4 nm and equivalent width 0.6 nm. Si IV (139.4, 140.3 nm) and He II (164.0 nm) are also seen with equivalent widths of 0.15 and 0.25 nm, respectively. Lyα absorption is apparent with full width likely > 10 000 km/s, which is suggestive of a massive accretor. No C IV or O V emission is detected, suggesting that the accreting material has been CNO-processed. The indicated evolutionary scenarios, coupled with the suggested 0.17-day orbital period (IAUC 7397), would make a primary of mass > 3 M_{⊙} unlikely. J. Stull, Z. Ioannou, and N. A. Webb obtained V -band photometry from the 0.8-m Stull Observatory telescope at Alfred University during Apr. 13.05–13.35. These data appear inconsistent with a 0.17-day period, suggesting instead a modulation at approximately twice this period. In this case, the 0.17day modulation might suggest nodal precession of a tilted disk with period approximately half the orbital period (http://xxx.lanl.gov/abs/astroph/?0004159). A 0.34-day period would remove the restriction to low-mass primaries referred to above. Further satellite observations are scheduled for Apr. 26–May 1 and May 4. Exact timings are given at http://phys-

ftp.open.ac.uk/pub/1118/satobs.”

O. Taranova and V. Shenavrin, Sternberg Astronomical Institute, report the following mean infrared magnitudes of XTE J1118+480 obtained during Apr. 12–15 with the 1.25-m telescope (+ InSb photometer) at the Crimean Station of the Sternberg Astronomical Institute: J = 12.4 ± 0.2; H = 11.9 ± 0.1; K = 10.9 ± 0.1; L = 9.2 ± 0.1.

SUPERNOVA 1999gv IN ANONYMOUS GALAXY Corrigendum. On IAUC 7405, line 5, for ‘broad B’ read ‘broad R’

2000 April 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green